OpenStudy (anonymous):

if f'(a)=0, then f has a local minimum of a local maximum at a.

OpenStudy (anonymous):

That is also false. Just because f'(a) = 0, doesnt guarantee a rel max/min.

OpenStudy (anonymous):

Example..take y = x^3 The derivative is 3x^2. set 3x^2 = 0, you get x = 0. But x = 0 is not a rel max NOT a rel min.

OpenStudy (anonymous):



OpenStudy (anonymous):

You can see from the graph of y = x^3, then x = 0 is not a rel max/min, although f ' (0) = 0.

OpenStudy (anonymous):

Makes sense?

OpenStudy (anonymous):

yes, if it's put in that form it's easier to understand.

OpenStudy (anonymous):

True. That's why I broke it down for you, so it is crystal clear.

OpenStudy (anonymous):

Thank you very much ^o^!

OpenStudy (anonymous):

welcome.